OpenStudy (anonymous):

find the cost of painting the sides of a water tank measuring 225 cm diameter and 375 in height at P20.00 per sq . m

OpenStudy (aripotta):

so it's a cylinder?

OpenStudy (anonymous):

I guess YES

OpenStudy (anonymous):

i got million!!! O__O

OpenStudy (aripotta):

LA cylinder = 2prh where p means pi 2p62.5(375) = 46875p = 147262.155637 cm^2 but now we have to convert cm to m

OpenStudy (anonymous):

\[2 \pi r \left( h+r \right)\]

OpenStudy (anonymous):

i got 35 per sq.m

OpenStudy (aripotta):

from our handy-dandy converter: 1472.6216 m^2

OpenStudy (aripotta):

then just multiply that by the cost: 1472.6216m x 20 = 29452.432 so, it would cost about P29452.43

OpenStudy (anonymous):

Just the sides, or end caps too?

OpenStudy (aripotta):

it said just the "sides"

OpenStudy (anonymous):

121.5 is the radius of 225 diameter

OpenStudy (aripotta):

it's 112.5. sheet.

OpenStudy (aripotta):

i really need some sleep lol.

OpenStudy (aripotta):

LA cylinder = 2prh where p means pi 2p112.5(375) = 84375p = 265071.880147 cm^2 converter: 2650.7188 m^2 cost: 2650.7188 x 20 = 53014.376 so it cost about P53014.38

OpenStudy (aripotta):

that should be right lol

OpenStudy (anonymous):

isn't 375 in inches?

OpenStudy (aripotta):

no, it says "in height"

OpenStudy (anonymous):

ah... took in to mean inches

OpenStudy (aripotta):

it did..............i will not be doing that again lol

OpenStudy (aripotta):

if it did*

OpenStudy (anonymous):

Gotta be careful... I got 530.14 When you convert square cm to square m.... there are 1000 square cm in one square m

OpenStudy (anonymous):

I converted all measurements to meters first, then calculated. Otherwise you have to deal with square units piling up.

OpenStudy (aripotta):

i used an online converter for all my conversions

OpenStudy (aripotta):

meh, whatever. i'm tired and should probably sleep lol good night :)

OpenStudy (anonymous):

265071 cm^2 is not 2650.71 m^2, it is 26.5071 m^2

OpenStudy (aripotta):

well, if the converter did it wrong, then whatever. :/

OpenStudy (anonymous):

k... goodnight